OpenStudy (anonymous):

Which is NOT a correct step when you solve the equation 5x - 3 = 8? A. 5^x=11 B.log511=x c.2^x=8 d. log11/log5=x

Directrix (directrix):

This looks like a linear equation the equation 5x - 3 = 8? But, the options look as if the equation has to do with logs. Please clarify. Thanks.

OpenStudy (anonymous):

Yes it is solving equations with logarithms

Directrix (directrix):

5x - 3 = 8 5x = 11 x = 11/5. Those steps appear to have nothing to do with the problem. I think you left something off the problem.

OpenStudy (anonymous):

The actual equation is written 5^x-3=8, my fault I thought i included it. So it actually 5 to the power of x.

Directrix (directrix):

5^x - 3 = 8 The 5^x and the 3 are not like terms so it would be incorrect to subtract them and get 2^x. c.2^x=8

OpenStudy (anonymous):

Thanks.